OpenStudy (clamin):

PLEASE HELP!! y=-3(x-4)^2+1 and pleasee show me how to graph it

OpenStudy (precal):

this is quadratic

OpenStudy (precal):

your vertex is (4, 1)

OpenStudy (clamin):

i know but what formula should i use??

OpenStudy (precal):

negative sign means reflection across x axis

OpenStudy (precal):

3 means vertical stretch your parent function is y=x^2 and is located on (0,0)

OpenStudy (hockeychick23):

y=-3(x-4)(x-4)+1 y=-3x^2+12x+12x-48+1 y= -3x^2+24x-47 x=4, y=1

OpenStudy (clamin):

how did u got 3x^2??

OpenStudy (hockeychick23):

y=-3(x-4)(x-4)+1 y= -3x^2+12x+12x-48+1 y= -3x^2+24x-47 x=4, y=1 Use this link to take you to a graphing calculator. It will graph the equation for you. https://www.desmos.com/calculator

OpenStudy (hockeychick23):

I got -3x^2 by distributing the -3 to the two x's

OpenStudy (clamin):

how'd you get 48??

OpenStudy (hockeychick23):

you need to distribute you take the -3 on the outside of the parentheses and multiply it by -4*-4. -3*-4*-4= -48